# The Value of Context: Human versus Black Box Evaluators

### v1

@@ 5.3 Sufficient residual uncertainty
Suppose Assumptions 6 and 7 hold. Then for every covariate vector $\mathbb{x}\in\{0,1\}^{\infty}$, the expected value of context vanishes to zero as $n$ grows large, i.e.,
@@ 6 Conclusion
Online versus offline learning. In our model, a key distinction between human and black box evaluation is that the human can adapt which covariates are acquired based on other properties of the agent, while the black box cannot. This is an appropriate comparison of human and black box evaluators as they currently stand: The black box algorithms used to make predictions about humans are usually supervised machine learning algorithms which are pre-trained on a large data set. But new black box algorithms, such as LLMs, blur this distinction, and future evaluations (e.g., medical diagnoses) may be conducted by black box systems with which the agent can communicate.
@@
In a change of notation relative to the main text, we subsequently use $\mathbb{X}_{n}$ to denote the agent’s covariate vector and $Y$ to denote the agent’s type (leaving $\mathbb{x}_{n}$ and $y$ to denote realizations of these random variables). Moreover, rather than supposing that $Y$ is deterministically related to $\mathbb{X}_{n}$ via a function $f$, let $(\mathbb{X}_{n},Y)\sim P^{n}$ where $P^{n}$ is unknown. We replace Assumption 1 with the following.
@@
That is, permuting the labels and/or values of the nonstandard covariates does not change the joint distribution of the conditional expectations of $y$. When $y$ is degenerate conditional on $\mathbb{x}_{n}$, Assumption 8 reduces to our previous Assumption 1. We will prove the following generalization of Theorem 3.1.
@@ A.2.4 Upper Bound for Expected Maximum of Gaussians
Finally by Berman (1964), which provides an upper bound for the expected maximum of independent Gaussian random variables, there exists a positive constant $C$ such that
$V^{N}_{n}\leq\frac{1}{2^{n-s-h_{n}}}C\sqrt{\log(K_{n})}\leq\frac{1}{2^{n(1-\alpha_{h})-s}}C\sqrt{n}$
@@
where the inner expectation converges to zero for every $\theta$ by Theorem A.2. By assumption that $u$ is Lipschitz continuous on a compact domain, there exist $\underline{u}$ and $\overline{u}$ such that $u(\hat{y},y)\in[\underline{u},\overline{u}]$ for all $(\hat{y},y)$. So $\mathbb{E}_{P^{n}\sim F_{\theta}}(v_{n}(P,\mathbb{x}_{n}))$ is pointwise bounded above by $\overline{u}-\underline{u}$, and we can apply the Dominated Convergence Theorem to conclude that $\lim_{n\rightarrow\infty}V(n,\mathbb{x})=0$ as desired.
@@
Throughout the proof we set $s=0$, $\mu=0$ and $\sigma^{2}=\mathbb{E}(Y_{i}^{2})=1$ without loss of generality. We’ll start by demonstrating Part (a). As before let $B_{n}\subseteq\{1,\dots,2^{n}\}$ index those $2^{n-b_{n}}$ covariate vectors that agree with the agent’s covariate vector for all covariates in $B$. Then the black box evaluator’s posterior expectation is the sample average
@@
for large enough $n$. We start by analyzing the first difference $\mathbb{E}[\phi(Z_{B}^{n})-\phi(0)]$. Using Taylor’s expansion we get
@@
$\mathbb{E}[\phi(Z_{B}^{n})-\phi(0)]\geq c_{1}\mathbb{E}[(Z_{B}^{n})^{2}]=\frac{c_{1}}{2^{(1-\alpha_{b})n}}$
@@
$\mathbb{E}\left[\max_{1\leq k\leq K_{n}}\phi(Z_{k})-\phi(0)]\leq c_{2}\mathbb{E}[\max\{|Z_{k}|\}\right]$
As a corollary of Lemma A.2 it follows that there exist $c_{3},c_{4}>0$ such that
$\mathbb{E}[\max\{|Z_{k}|\}]\leq c_{3}\frac{1}{2^{(1-\alpha_{h})n}}\sqrt{\log(c_{4}K_{n})}$
@@
$\mathbb{E}\left[\max_{1\leq k\leq K_{n}}\phi(Z_{k})-\phi(0)\right]\leq c_{2}c_{3}\frac{1}{2^{(1-\alpha_{h})n}}\sqrt{\log(c_{4}K_{n})}$
@@
$\Delta(n)\geq c_{1}\frac{1}{2^{(1-\alpha_{b})n}}-c_{2}c_{3}\frac{1}{2^{(1-\alpha_{h})n}}\sqrt{\log(c_{4}K_{n})}$
The RHS is asymptotically positive if and only if
@@
since $\sqrt{\log(c_{4}K_{n})}$ has sub-exponential but non-constant asymptotics. This condition is satisfied if and only if $\alpha_{b}>\alpha_{h}$.
Part (b) follows by identical arguments: Since $-\phi$ is convex, the above arguments apply to show that
@@
$\geq-c_{2}c_{3}\frac{1}{2^{(1-\alpha_{h})n}}\sqrt{\log(c_{4}K_{n})}$
for some $c_{2},c_{3},c_{4}>0$. The desired conclusion follows.
@@ B.3 Result for Mixed Strategy Equilibria
to be the most-preferred achievable posterior expectation, which is unique by assumption of genericity of the equilibrium. Since $Z_{(1)}^{*}$ is the best attainable posterior expectation, an agent achieves $Z_{(1)}^{*}$ in equilibrium if and only if it is feasible. (Otherwise, the agent can profitably deviate to the feasible disclosure that induces this posterior expectation.)
@@
where $p_{\mathcal{X}^{*}}$ is the ex-ante probability that the agent’s covariate vector belongs to $\mathcal{X}^{*}$, and $\mathbb{E}[Y|X\notin\mathcal{X}^{*}]$ is the expectation of the agent’s type given that his covariate vector does not belong to $\mathcal{X}^{*}$. Here we utilize the fact that the evaluator’s posterior expectation is constant at $Z_{(1)}^{*}$ across all agents with covariate vectors in $\mathcal{X}^{*}$. [^30] In general this does not have to be the case. We rule this out in the definition of the equilibrium. Now, consider the following alternative “strategy” $\sigma_{0}$, which relaxes the feasibility constraint: For any $\mathbb{x}\in\mathcal{X}\setminus\mathcal{X}^{*}$ let $\sigma_{0}(\textbf{x})\equiv\sigma^{*}(\textbf{x})$, i.e., the disclosures are the same as in the original equilibrium. Further choose some arbitrary disclosure $d_{0}\in\mathcal{D}(\mathcal{X}^{*})$ and let $\sigma_{0}(\textbf{x})=d_{0}$ for all $\mathbb{x}\in\mathcal{X}^{*}$. The Receiver’s posterior expectation following observation of disclosure $d_{0}$ is
@@
We again continue in the general setting outlined in the proof of Theorem A.1, and adopt the conventions that $\mathbb{E}(Y)=\mu$ while $\text{Var}(Y)=1$. Recalling that $r_{n}$ is the number of relevant covariates, there are $2^{r_{n}}$ distinct expected conditional types, which we can enumerate as $Y_{1},\dots,Y_{2^{r_{n}}}$. If disclosure $k$ involves disclosing $k_{r}$ relevant covariates, then there is a set $S_{k}$ of size $2^{r_{n}-k_{r}}$ such that the evaluator’s posterior expectation can be written
@@
B.5 Proof of Proposition 5.2
@@
By Assumption 6, inequality B.4 simplifies to

### v2

@@ 5.4 Sufficient residual uncertainty
Suppose Assumptions 9 and 10 hold. Then for every covariate vector $\mathbb{x}\in\{0,1\}^{\infty}$, the expected value of context vanishes to zero as $n$ grows large, i.e.,
@@ 6 Conclusion
Online versus offline learning. In our model, a key distinction between human and black box evaluation is that the human can adapt which covariates are acquired based on other properties of the agent, while the black box cannot. This is an appropriate comparison of human and black box evaluators as they currently stand: The black box algorithms used to make predictions about humans are usually supervised machine learning algorithms which are pre-trained on a large data set. But new black box algorithms, such as large language models, blur this distinction, and future evaluations (e.g., medical diagnoses) may be conducted by black box systems with which the agent can communicate.
@@
In a change of notation relative to the main text, we subsequently use $\mathbb{X}_{n}$ to denote the agent’s covariate vector and $Y$ to denote the agent’s type (leaving $\mathbb{x}_{n}$ and $y$ to denote realizations of these random variables). Moreover, rather than supposing that $Y$ is deterministically related to $\mathbb{X}_{n}$ via a function $f$, let $(\mathbb{X}_{n},Y)\sim P^{n}$ where $P^{n}$ is unknown. We replace Assumptions 1 and 2 with the following weaker assumption.
@@
That is, permuting the labels and/or values of the nonstandard covariates does not change the joint distribution of the conditional expectations of $y$. When $y$ is degenerate conditional on $\mathbb{x}_{n}$, Assumption 11 reduces to our previous two assumptions. We will prove the following generalization of Theorem 3.1.
@@ A.2.4 Upper Bound for Expected Maximum of Gaussians
Finally by Berman (1964), which provides an upper bound for the expected maximum of independent Gaussian random variables
$V^{N}_{n}\leq\frac{1}{2^{n-s-h_{n}}}\cdot 2\sqrt{\log(K_{n})}\leq\frac{1}{2^{n(1-\alpha_{h})-s}}\cdot 2\sqrt{n}$
@@
where the inner expectation converges to zero for every $\theta$ by Theorem A.2. By assumption that $u$ is Lipschitz continuous on a compact domain, there exist $\underline{u}$ and $\overline{u}$ such that $u(\hat{y},y)\in[\underline{u},\overline{u}]$ for all $(\hat{y},y)$. So $\mathbb{E}_{P^{n}\sim F_{\theta}}(v_{n}(P,\mathbb{x}_{n}))$ is pointwise bounded above by $\overline{u}-\underline{u}$, and the Dominated Convergence Theorem implies $\lim_{n\rightarrow\infty}V(n,\mathbb{x})=0$, as desired.
@@
Throughout the proof we set $s=0$, $\mu=0$ and $\sigma^{2}=\mathbb{E}(Y_{i}^{2})=1$ without loss of generality. We’ll start by demonstrating that the stated results hold asymptotically (i.e., for large enough $n$) and subsequently prove that the bound in (3.4) is sufficient.
(a) As before let $B_{n}\subseteq\{1,\dots,2^{n}\}$ index those $2^{n-b_{n}}$ covariate vectors that agree with the agent’s covariate vector for all covariates in $B$. Then the black box evaluator’s posterior expectation is the sample average
@@
for large enough $n$.
We start by analyzing the first difference $\mathbb{E}[\phi(Z_{B}^{n})-\phi(0)]$. Using Taylor’s expansion we get
@@
$\mathbb{E}[\phi(Z_{B}^{n})-\phi(0)]\geq c_{1}\mathbb{E}[(Z_{B}^{n})^{2}]=\frac{c_{1}}{2^{(1-\alpha_{b})n}}$ (A.14)
@@
$\mathbb{E}\left[\max_{1\leq k\leq K_{n}}\phi(Z_{k})-\phi(0)\right]\leq c_{2}\mathbb{E}\left[\max\{|Z_{k}|\}\right]$
From our proof of Proposition A.2 it follows that
$\mathbb{E}[\max\{|Z_{k}|\}]\leq\frac{1}{2^{(1-\alpha_{h})n-1}}\sqrt{\log(K_{n})}$
@@
$\mathbb{E}\left[\max_{1\leq k\leq K_{n}}\phi(Z_{k})-\phi(0)\right]\leq c_{2}\frac{1}{2^{(1-\alpha_{h})n-1}}\sqrt{\log(K_{n})}$
@@
$\Delta(n)\geq c_{1}\frac{1}{2^{(1-\alpha_{b})n}}-2c_{2}\frac{1}{2^{(1-\alpha_{h})n}}\sqrt{\log(K_{n})}$
The RHS is positive for all large $n$ if and only if
@@
since $\sqrt{\log(K_{n})}$ has sub-exponential but non-constant asymptotics. This condition is satisfied if and only if $\alpha_{b}>\alpha_{h}$.
(b) Since $-\phi$ is convex, the above arguments apply to show that
@@
$\geq-c_{2}\frac{1}{2^{(1-\alpha_{h})n-1}}\sqrt{\log(K_{n})}$
for some $c_{2}>0$. The desired conclusion follows.
Finally, we show that the bound in (3.4) is sufficient for the comparison in part $(a)$ of the result (with identical computations applying to part $(b)$). Suppose $\phi(.)$ is strictly convex and denote $C=\frac{2c_{2}}{c_{1}}$, where $c_{1},c_{2}>0$ are the constants used above, respectively reflecting $\phi$’s lowest degree of convexity ($c_{1}=\inf_{y\in[-\overline{y},\overline{y}]}|\phi^{\prime\prime}(y)|$) and largest growth rate ($c_{2}=\sup_{y\in[-\overline{y},\overline{y}]}|\phi^{\prime}(y)|$). Following the proof of part (a), the Black Box is preferred if
$2^{(\alpha_{b}-\alpha_{h})n}>C\sqrt{\log(K_{n})}$
Since $K_{n}\leq 2^{n}$, this inequality is satisfied if
$(\alpha_{b}-\alpha_{r})n-\frac{1}{2}\log_{2}(n)>\log_{2}(C)$
The above inequality implicitly defines a threshold on the sufficient number of covariates $N(C)$, exceeding which the Black Box is preferred.
@@ B.3 Result for Mixed Strategy Equilibria
to be the most-preferred achievable posterior expectation, which is unique by assumption of genericity of the equilibrium.
Since $Z_{(1)}^{*}$ is the best attainable posterior expectation, an agent achieves $Z_{(1)}^{*}$ in equilibrium if and only if it is feasible. (Otherwise, the agent can profitably deviate to the feasible disclosure that induces this posterior expectation.)
@@
where $p_{\mathcal{X}^{*}}$ is the ex-ante probability that the agent’s covariate vector belongs to $\mathcal{X}^{*}$, and $\mathbb{E}[Y|X\notin\mathcal{X}^{*}]$ is the expectation of the agent’s type given that his covariate vector does not belong to $\mathcal{X}^{*}$. Here we utilize the fact that the evaluator’s posterior expectation is constant at $Z_{(1)}^{*}$ across all agents with covariate vectors in $\mathcal{X}^{*}$. [^30] In general this does not have to be the case. We rule this out in the definition of the equilibrium.
Now, consider the following alternative “strategy” $\sigma_{0}$, which relaxes the feasibility constraint: For any $\mathbb{x}\in\mathcal{X}\setminus\mathcal{X}^{*}$ let $\sigma_{0}(\textbf{x})\equiv\sigma^{*}(\textbf{x})$, i.e., the disclosures are the same as in the original equilibrium. Further choose some arbitrary disclosure $d_{0}\in\mathcal{D}(\mathcal{X}^{*})$ and let $\sigma_{0}(\textbf{x})=d_{0}$ for all $\mathbb{x}\in\mathcal{X}^{*}$. The Receiver’s posterior expectation following observation of disclosure $d_{0}$ is
@@
We again continue in the general setting outlined in the proof of Theorem A.1, and adopt the conventions that $\mathbb{E}(Y)=\mu$ while $\text{Var}(Y)=1$. We prove the result for a weakening of Assumptions 5 and 6 to the following.
Assumption 13.
Recalling that $r_{n}$ is the number of relevant covariates, there are $2^{r_{n}}$ distinct expected conditional types, which we can enumerate as $Y_{1},\dots,Y_{2^{r_{n}}}$. If disclosure $k$ involves disclosing $k_{r}$ relevant covariates, then there is a set $S_{k}$ of size $2^{r_{n}-k_{r}}$ such that the evaluator’s posterior expectation can be written
@@
B.5 Supporting Materials for Section 5.3
We show here that both main results generalize to the setting with noise. Instead of repeating the proof step by step, we emphasize what changes must be made in order for the proofs to translate. To be consistent with our previous notation, we again operate with $\tilde{y}_{i,n}=Y_{i}+\varepsilon_{n}$ where $i$ enumerates covariate vectors. In all subsequent notation a tilde will indicate an object that includes noise $\varepsilon_{n}$, and objects without one are the same as in the main text. Observe that after adding the noise term $\varepsilon_{n}$, the key objects from the proofs transform in the following way: $Z_{k}$ is replaced by $\tilde{Z}_{k}=Z_{k}+\varepsilon_{n}$, and $\mathbb{E}[\max_{k}Z_{k}]$ is replaced by $\mathbb{E}[\max_{k}\tilde{Z}_{k}]$. Since $\mathbb{E}[\max_{k}\tilde{Z}_{k}]=\mathbb{E}[\max_{k}Z_{k}]$, the proof of Theorem 3.1 translates directly.
To demonstrate Theorem 3.2, we will show how to adjust the proof of part (a) with part (b) following analogously. The first change is that (A.14) becomes
$\mathbb{E}[\phi(\tilde{Z}_{B}^{n})-\phi(0)]\geq c_{1}\mathbb{E}[(Z_{B}^{n})^{2}]-c_{1}\sigma_{\varepsilon,n}^{2}$
The inequality in (A.15) is also modified to
$\phi(\tilde{Z}_{k})-\phi(0)\leq c_{2}|Z_{k}|+c_{2}|\varepsilon_{n}|$
for every disclosure $k$. Thus we obtain
$\Delta(n)\geq c_{1}\frac{1}{2^{(1-\alpha_{b})n}}-2c_{2}\frac{1}{2^{(1-\alpha_{h})n}}\sqrt{\log(K_{n})}+c_{2}\mathbb{E}[|\varepsilon_{n}|]-c_{1}\sigma_{\varepsilon,n}^{2}$
We will show that the ratio $\frac{\mathbb{E}[|\varepsilon_{n}|]}{\sigma_{\varepsilon,n}^{2}}$ grows arbitrary large with $n$, thus asymptotically exceeding $\frac{c_{1}}{c_{2}}$. Fixing some $d>0$ and applying Markov inequality, we obtain
$\frac{\mathbb{E}[|\varepsilon_{n}|]}{\sigma_{\varepsilon,n}^{2}}\geq d\cdot\mathbb{P}\left(\frac{|\varepsilon_{n}|}{\sqrt{\sigma_{\varepsilon,n}^{2}}}\geq d\sqrt{\sigma_{\varepsilon,n}^{2}}\right)$ (B.4)
If we denote the CDF of $\frac{\varepsilon_{n}}{\sqrt{\sigma_{\varepsilon,n}^{2}}}$ as $G_{n}$, the RHS of the above inequality can be rewritten as $d\cdot\left(1+G_{n}(-d\sqrt{\sigma_{\varepsilon,n}^{2}}))-G_{n}(d\sqrt{\sigma_{\varepsilon,n}^{2}})\right)$. As $n$ grows large, the term in brackets tends to $1-2g_{n}(0)d\sqrt{\sigma_{\varepsilon,n}^{2}}+o\left(\sqrt{\sigma_{\varepsilon,n}^{2}}\right)$. We will omit the $o(\cdot)$ term until the end of the proof.
Fix an arbitrary $\delta>0$ and let $d=\frac{c_{1}}{c_{2}}+2\delta$. Further, fix $N$ such that $\sqrt{\sigma_{\varepsilon,n}^{2}}\leq\frac{1}{2gd}(1-\frac{c_{1}+c_{2}\delta}{c_{1}+2c_{2}\delta})$, where $g=\max_{n}g_{n}(0)<\infty$. Then since $\sigma_{\varepsilon,n}^{2}$ is decreasing and $g_{n}(0)\leq g$ we have that for all $n\geq N$
$2g_{n}(0)d\sqrt{\sigma_{\varepsilon,n}^{2}}\leq 2gd\frac{1}{2gd}\left(1-\frac{c_{1}+c_{2}\delta}{c_{1}+2c_{2}\delta}\right)$
Combining this inequality with (B.4) we get
$\frac{\mathbb{E}[|\varepsilon_{n}|]}{\sigma_{\varepsilon,n}^{2}}\geq\left(\frac{c_{1}}{c_{2}}+2\delta\right)\left(1-\left(1-\frac{c_{1}+c_{2}\delta}{c_{1}+2c_{2}\delta}\right)\right)+o\left(\sqrt{\sigma_{\varepsilon,n}^{2}}\right)=\left(\frac{c_{1}}{c_{2}}+\delta\right)+o\left(\sqrt{\sigma_{\varepsilon,n}^{2}}\right)$
for all $n\geq N$. Since $\delta$ is an arbitrary positive number, this concludes the proof.
B.6 Proof of Proposition 5.2
@@
By Assumption 9, inequality B.5 simplifies to